OpenStudy (hellomynameiskam):

Solve the equation. - 3 2/5 + b = 8 1/5 A. b = 11 3/5 B. b= 4 4/5 C. b 5 3/5 D. b = 11 1/5 @kendricklamar2014

OpenStudy (xhawaiiangirlx):

A or D

OpenStudy (hellomynameiskam):

im not sure. @KendrickLamar2014

OpenStudy (abdullahm):

First, convert the fractions into improper fractions. Do you know how to do that?

OpenStudy (hellomynameiskam):

only a lil. my mind went blank after the week i learned it.

OpenStudy (abdullahm):

\(\Large a\frac{b}{c} = \frac{(a~\cdot~ c)~ + ~b}{c}\)

OpenStudy (hellomynameiskam):

oh okay

OpenStudy (hellomynameiskam):

i got - 17/5 and 46/5

OpenStudy (abdullahm):

Check the second improper fraction again

OpenStudy (hellomynameiskam):

41/5

OpenStudy (abdullahm):

correct

OpenStudy (abdullahm):

We now have: \(\Large- \frac{17}{5} + b = \frac{41}{5}\) Rearrange the terms \(\Large b - \frac{17}{5}= \frac{41}{5}\) Add 17/5 on both sides \(\Large b = \frac{41}{5}+\frac{17}{5}\) Can you simplify that more? :)

OpenStudy (hellomynameiskam):

I would have to say yes. Even though i am not sure i would have to say yes

OpenStudy (abdullahm):

Then simplify it :P

OpenStudy (abdullahm):

Hint: \(\Large \frac{a}{c} + \frac{b}{c} = \frac{a~+~b}{c}\)

OpenStudy (hellomynameiskam):

i got 58/ 5 so that has to no to simplifying right?

OpenStudy (abdullahm):

We can convert 58/5 into a mixed number, or we can leave it as an improper fraction. :)

OpenStudy (abdullahm):

The answer choices are all in mixed number, so let's go ahead and make 58/5 into a mixed number. :)

OpenStudy (hellomynameiskam):

58 divided by 5 is 11.6

OpenStudy (abdullahm):

Which answer choice do you think it is? :)

OpenStudy (hellomynameiskam):

d?

OpenStudy (abdullahm):

Nope. D is equal to 11.2

OpenStudy (hellomynameiskam):

ohh so its A

OpenStudy (abdullahm):

Correct! :)

OpenStudy (hellomynameiskam):

i have 2 more. can u help?